Circular No. 7209

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cq IN UGC 11268 M. Modjaz and W. D. Li, University of California at Berkeley, on behalf

of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), report the discovery of an apparent supernova in UGC 11268. The supernova was discovered on an unfiltered images taken on June 25.4 UT (mag ∼ 15.8), and confirmed by earlier observations on June 22.4 (mag ∼ 15.8) and 19.4 (mag ∼ 15.9). The new object is located at α = 18^{h}32^{m}07^{s}.10, δ = +37^{o}36′44′′.3 (equinox 2000.0), which is 1′′.5 east and 4′′.1 south of the nucleus of UGC 11268. A KAIT image of the same field on June 15.4 UT (limiting mag ∼ 19.0) showed nothing at the position of the supernova.

NOVA IN NGC 6822 J. Y. Wei, D. W. Xu, Q. Y. Qiao, Y. L. Qiu and J. Y. Hu, Beijing

Astronomical Observatory (BAO), report the independent discovery by the BAO Supernova Survey of the object reported on IAUC 7208. The observations with the 0.6-m telescope at Xinglong showed the apparent nova on June 23.69 and 24.72 UT at unfiltered mag ∼ 18.

V382 VELORUM Additional photometry by A. C. Gilmore with the 0.6-m f/16 reflector

at Mt. John University Observatory (same reference star data as on IAUC 7196): June 9.488 UT, V = 6.41, U − B = −0.92, B − V = −0.07, V − R = +1.34, V − I = +0.94, airmass = 1.63; 10.368, 6.46, −0.93, −0.09, +1.34, +0.93, 1.12; 18.366, 6.97, −0.95, −0.17, +1.42, +0.92, 1.17; 19.457, 7.09, −0.86, −0.18, +1.46, +0.95, 1.60. Standard deviations are generally < 0.01 mag, except in U − B, where it is generally 0.05 mag.

Visual magnitude estimates by W. C. de Souza, São Paulo, Brazil: June 21.90 UT, 7.3; 23.01, 7.3; 24.91, 7.3; 25.90, 7.3.

COMET C/1999 H1 (LEE) Total visual magnitude estimates: June 10.28 UT, 6.7 (S. T. Rae,

Palmerston North, New Zealand, 10 × 50 binoculars); 14.29, 6.7 (N. Biver, Oahu, HI, 7×50 binoculars); 17.35, 6.5 (D. A. J. Seargent, Cowra, N.S.W., Australia, 25 × 100 binoculars); 19.36, 6.5 (C. E. Drescher, Warrill View, Qld., Australia, 20 × 80 binoculars; 21.89, 6.9 (W. C. de Souza, São Paulo, Brazil, 11 × 80 binoculars).

1999 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden